What can you say about the upper bound of E(x) for numbers of the form 3^(2n) + 1?


1
Expert's answer
2021-12-14T13:30:27-0500

Consider the set

"\begin{aligned}

&E(x)=\left\{3^{3 n}+1: n \in N 1\right\} . \\

&E(x) \neq \phi . \text { since } 3^{2}+1 \in E(x) . \\

&E(x) \subseteq \mathbb{R} .

\end{aligned}"

Now if c is the least upper of E(x).

"c=\ln _{b}((-(x))."

"\forall \epsilon>0, \quad c-\epsilon<3^{2 n+1}, \quad[\because c-\epsilon, is not

* n \in N. lub \ of \ E(x)]"

Take

"c-2 \cdot 3^{2 n}<3^{2 n}+1 \quad \forall n \in N \text {. }"

"\begin{aligned}

&c<3^{2 n}+8 \cdot 3^{2 n}+1, \quad \forall n \in N . \\

&c<3^{2 n}(1+8)+1 . \\

&c<3^{2 n} \cdot 3^{2}+1 \\

&c<3^{2(n+1)}+1 .

\end{aligned}"

c is not lub of E(x)

Thus E(x) has no upper bound.